OpenStudy (anonymous):

Consider the equation: xsec(y)=y-1 A. compute dy/dx B compute the slope of this curve at the point (-1,0)

hartnn (hartnn):

tried part a ?

hartnn (hartnn):

u know implicit differentiation ?

OpenStudy (anonymous):

i know that every time i take a derivative with a y in it ill get dy/dx times it as well. Then i need to group all with dy/dx on one side and the other side group all without them. Then divide to get dy/dx= the answer.

hartnn (hartnn):

thats almost correct! so whats your attempt in this Q ?

OpenStudy (anonymous):

on the left side i need to take the derivative of xsec(y) so starting with the product rule yes?

hartnn (hartnn):

yup, go on...i'll stop u when u do any error...

OpenStudy (anonymous):

or you could take dx/dy and invert it at the end

hartnn (hartnn):

you'll need a product rule anyways...

OpenStudy (anonymous):

true , but it'll end up with the same answer either way

OpenStudy (anonymous):

Alright well taking x'sec(y) 1sec(y)+ x*sec(y)' so you get sec(y)+(dy/dx)xsec(y)tan(y) right?

hartnn (hartnn):

yes, and on right ?

OpenStudy (anonymous):

the derivative of y-1 is just 1(dy/dx) right?

OpenStudy (anonymous):

yes .. now group the dy/dx terms to one side of the equation

hartnn (hartnn):

doing good job! go on..

OpenStudy (anonymous):

so sec(y)=(dy/dx)-(dy/dx)sec(y)tan(y) factoring out dy/dx you get sec(y)=dy/dx(1-xsec(y)tan(y)) so dy/dx=sec(y)/(1-xsec(y)tank(y))?

hartnn (hartnn):

seems correct.

OpenStudy (anonymous):

for B , slope is dy/dx at the specifies point

OpenStudy (anonymous):

specified*

hartnn (hartnn):

you just need to put x=-1, y=0 in dy/dx you got.

OpenStudy (anonymous):

So the slope is 1

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thank you very much for your help

hartnn (hartnn):

welcome but you did all the work ! :)